A PROJECTILE IS PROJECTED WITH A VELOCITY (30i+40j)m/s . g=10m/s^2. Then its velocity at t= 1 second is
 

Grade:12th pass

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Vx = 30

Vy = 40

As no horizontal force is present no change in horizontal component

Vx after 1 sec = 30 m/s

Vy = 40 - 10(1)
Vy = 30 m/s

V = ( 30i +30j) m/s

Regards
Khimraj
4 years ago
Vx will remain constant
Vy = 40 – 10*1 = 30
V = Vx + Vy
V = ( 30i +30j) m/s
….............................................................................................................